What will be the average density for a sample of metal with a mass of 12.6 g and volume calculations of 5.5 mL and 8.6 cm3 ? Omit units in the answer and report answer to one decimal place
Answer #1

density is ratio between mass and volume. Average density is the ratio between sum of mass and sum of volume.

D=M/V

D=(12.6+12.6)/(5.5+8.6)

= 25.2/14.1

= 1.8